OpenStudy (anonymous):

If sin A = 15/17, What is the measurement of ∠Α and ∠B ?

jimthompson5910 (jim_thompson5910):

sin A = 15/17 arcsin(sin A) = arcsin(15/17) ... apply the arcsine function to both sides A = 61.927513 ... use a calculator (this is approximate)

jimthompson5910 (jim_thompson5910):

I'll let you compute angle B

OpenStudy (anonymous):

Ok how would i compute this without a TI Calculator?

jimthompson5910 (jim_thompson5910):

with a table is your next best bet

OpenStudy (anonymous):

Thats my issue with understanding these questions

jimthompson5910 (jim_thompson5910):

you can't do something like this by hand

jimthompson5910 (jim_thompson5910):

I'd use a calculator if you can

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

keep in mind that arcsin is the same as \(\large \sin^{-1}\) or inverse sine

OpenStudy (anonymous):

ok, that i am not familiar with. I just got help with cos and tan.

jimthompson5910 (jim_thompson5910):

if you have a TI calculator, the inverse sine is found by hitting 2nd, then the sine button

OpenStudy (anonymous):

oh i dont have one

OpenStudy (anonymous):

using a regular calculator

jimthompson5910 (jim_thompson5910):

ok you can use google as a calculator type "arcsin(15/17) in degrees" without quotes into google to get https://www.google.com/search?q=arcsin%2815%2F17%29+in+degrees&ie=UTF-8

OpenStudy (anonymous):

Oh ok thank you.

OpenStudy (anonymous):

Let me work this out i will come back and see if i did it right

OpenStudy (anonymous):

to get the measure of angle B = arcos(15/17) = arcsin(8/15)

jimthompson5910 (jim_thompson5910):

it's arccos and not arcos

OpenStudy (anonymous):

oh sorry

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

but am i correct for 8/15

jimthompson5910 (jim_thompson5910):

and arcos(15/17) = arcsin(8/15) is simply not true at all

OpenStudy (anonymous):

Oh ok i will try again

jimthompson5910 (jim_thompson5910):

if you meant arccos(15/17) = arcsin(8/17), then you are correct

OpenStudy (anonymous):

for my answer i got 25.20

jimthompson5910 (jim_thompson5910):

for which one again?

OpenStudy (anonymous):

for angle B

jimthompson5910 (jim_thompson5910):

close, but it's a bit smaller than it should be

jimthompson5910 (jim_thompson5910):

hint: A+B = 90

OpenStudy (anonymous):

oh dang it..

jimthompson5910 (jim_thompson5910):

A and B are complementary angles

OpenStudy (anonymous):

Oh i see where i went wrong. Let me try again.

OpenStudy (anonymous):

Ok i think i got it now, The measure of angle B=28.07

jimthompson5910 (jim_thompson5910):

yes approximately

OpenStudy (anonymous):

Thank you.

jimthompson5910 (jim_thompson5910):

you're welcome